# Mapping Cone Connections and their Yang-Mills Functional

### v2

@@ 1 Introduction
In this paper, we study a Yang-Mills type functional that involves a pair $(A,B)$, with $A\in\Omega^{1}(M,Ad\,P)$ being the connection one-form and $B\in\Omega^{0}(M,Ad\,P)$ a section of the associated adjoint bundle $Ad\,P$, and also a two-form $\zeta\in\Omega^{2}(M)$ that is $d$-closed. We shall call this functional the cone Yang-Mills functional and it is defined by
@@
which take values on the associated adjoint bundle $Ad\,P$. In this context, the differential $d_{C}$ must also be twisted by a cone connection one-form $\mathcal{A}\in\operatorname{Cone}^{1}(\zeta)(M,Ad\,P)$
@@
where $\mathcal{A}=A+\theta\wedge B\,$, with $A\in\Omega^{1}(M,Ad\,P)$ and $B\in\Omega^{0}(M,Ad\,P)$. The cone curvature then takes the form
@@
Acknowledgements.  We would like to thank Alberto Cattaneo, David Clausen, Bernard Julia, Claude LeBrun, Si Li, Jianfeng Lin, Xiang Tang, Daniel Waldram, Jiaping Wang, Richard Wentworth, and Dan Xie for helpful comments and suggestions. We would like to acknowledge the support of the Simons Collaboration Grant No. 636284 for the first author and the support of the National Key Research and Development Program of China No. 2020YFA0713000 for the second author.
@@ 2.1 Preliminaries
Let $M$ be a smooth manifold and $P$ be a principal bundle over $M$. We consider the associated adjoint bundle $Ad\,P$ equipped with an inner product that is invariant under the adjoint action. Hence, for arbitrary $x,y,z\in\Omega^{*}(M,Ad\,P)$,
$\langle[x,y],z\rangle=\langle x,[y,z]\rangle.$ (2.1)
@@
In this paper, we take the inner product on $\Omega^{*}(M,Ad\,P)$ to be the standard one induced from the Riemannian metric $g$ and the Killing form of the Lie group $G$. Specifically,
where the constant $c_{G}=N-2$ if $G=SO(N)$. With (2.2) and (2.3), we can also express $\langle-,-\rangle_{\mathcal{C}}$ as an integral by introducing the Hodge star operator for cone forms
@@
Having defined the inner product on $\operatorname{Cone}^{*}(\zeta)(M,Ad\,P)$, we write down the Yang-Mills functional associated with the cone connection. Given a connection $A\in\Omega^{1}(M,Ad\,P)$ and a section $B\in\Omega^{0}(M,Ad\,P)$, we write the cone connection as $\mathcal{A}=A+\theta B\in\operatorname{Cone}^{1}(\zeta)(M,Ad\,P)$. Locally, if we write the usual covariant derivative as $d_{A}=d+A$, then the cone covariant derivative is given by $D_{\mathcal{C}}=d+\mathcal{A}=d+(A+\theta B)$ and the cone curvature takes the form
@@ 2.3 Comparison with Yang-Mills solutions
Let $\Sigma$ be a Riemann surface and $A^{\prime}\in\Omega^{1}(\Sigma,Ad\,P)$ be a non-flat Yang-Mills connection on $\Sigma$. Let $\zeta$ be the volume form of $\Sigma$ normalized such that the total volume of $\Sigma$ is one. Define $M$ to be the three-dimensional circle bundle $\pi:M\to\Sigma$ with Euler class given by $\zeta$. Since $A^{\prime}$ is Yang-Mills, we can write its curvature as $F_{A^{\prime}}=\zeta\Phi$ such that $d_{A^{\prime}}\Phi=0$. For simplicity, we will also use $\zeta$, $A^{\prime}$ and $\Phi$ to denote their pullbacks on $M$.
@@
With $\zeta$ being a non-degenerate 2-form, (2.21) gives the condition $dB=0$.
@@ 2.4 Uniqueness of cone Yang-Mills solutions for a fixed connection
However, it is interesting to ask that given a cone Yang-Mills solution $(A,B)$, how many different $B$’s with $A$ fixed would also be a cone Yang-Mills solution? In the case of $\zeta=0$ and $M$ closed, if $(A,B)$ is a cone Yang-Mills solution, then so are all $(A,B+\Phi)$ with $d_{A}\Phi=0$. For $\zeta\neq 0$, we are able to obtain results in two special cases. First, for two dimensions and suppose $A$ is a Yang-Mills connection, then we have the following:
@@
Let $\dim M=2n$. With respect to a compatible metric, $*\,\zeta=*\,\omega=\frac{1}{(n-1)!}\omega^{n-1}$, and $\zeta^{*}\zeta=n$
@@ 3.2 Three-dimensional solutions from duality relations
conformally scaled by the $e^{2\phi}$ factor in order match the desired metric form of $ds^{2}_{X}=ds^{2}_{M}+\theta^{2}$ as in (1.5). We note that the four-dimensional Yang-Mills anti-self-dual/self-dual curvature equation is invariant under an overall conformal scaling of the metric. Therefore, for the cone pair $(A,B)$ on $M=\mathbb{R}^{3}-\{0\}$, we can just read off from (3.11)-(3.12) by setting $\mathcal{A}^{r}{}_{s}=(A+\theta B)^{r}{}_{s}=\omega^{r}{}_{s}\,$. We find
@@
We can consider dimensionally reducing the three-dimensional duality condition $*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\mp\,\mathcal{F}_{\mathcal{A}}$ over another circle down to a two-dimensional manifold $N$. For the metric, we will assume $ds_{M}^{2}=ds_{N}^{2}+\chi^{2}$ where $\chi$ is the global connection one-form (i.e. the global angular form) of the circle bundle $M$ over $N$. The dimensionally reduced three-dimensional local connection can then be expressed as $A+\chi b$ where $A\in\Omega^{1}(N,Ad\,P)$ and $b\in\Omega^{0}(N,Ad\,P)$. The three-dimensional cone curvature then takes the form
@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
To simplify our consideration, we will additionally assume that $\zeta$ is a non-degenerate two-form. Being $d$-closed and non-degenerate, $\zeta$ is then a symplectic structure and $M$ must be even-dimensional. Additionally, $\zeta$ being non-degenerate leads to two simplifications in considering cone-flat pairs $(A,B)$. First, non-degeneracy ensures that $\zeta$ nowhere vanishes, and hence, the first cone-flat equation $F_{A}=-\zeta B$, implies $B$ is determined by the connection $A$. Second, the second cone-flat equation $d_{A}B=0$ is automatically satisfied and redundant when $\dim M\geq 4$. For the first cone-flat equation together with the Bianchi identity imply

### v3

@@ 1 Introduction
In this paper, we study a Yang-Mills type functional that involves a pair $(A,B)$, with $A$ being a connection form on the associated adjoint bundle $Ad\,P$, a section $B\in\Omega^{0}(M,Ad\,P)$, and also a two-form $\zeta\in\Omega^{2}(M)$ that is $d$-closed. We shall call this functional the cone Yang-Mills functional and it is defined by
@@
which take values on the associated adjoint bundle $Ad\,P$. In this context, the differential $d_{C}$ must also be twisted by a cone connection form $\mathcal{A}$
@@
where $\mathcal{A}=A+\theta\wedge B\,$, with $A$ being the connection form on $M$ and $B\in\Omega^{0}(M,Ad\,P)$. The cone curvature then takes the form
@@
Acknowledgements.  We would like to thank Alberto Cattaneo, David Clausen, Jonathan Delgado, Bernard Julia, Claude LeBrun, Si Li, Jianfeng Lin, Xiang Tang, Daniel Waldram, Jiaping Wang, Richard Wentworth, and Dan Xie for helpful comments and suggestions. We would like to acknowledge the support of the Simons Collaboration Grant No. 636284 for the first author and the support of the National Key Research and Development Program of China No. 2020YFA0713000 for the second author.
@@ 2.1 Preliminaries
Let $M$ be a smooth manifold and $P$ be a principal bundle over $M$. We consider the associated adjoint bundle $Ad\,P$ equipped with an inner product that is invariant under the adjoint action. For the inner product on $\Omega^{*}(M,Ad\,P)$, we use the standard one induced from the Riemannian metric $g$ and the Killing form of the Lie group $G$, i.e. for $\eta_{1},\eta_{2}\in\Omega^{*}(M,Ad\,P)$,
where the constant $c_{G}=-(N-2)$ if $G=SO(N)$. As will be useful later, we note here for $\eta\in\Omega^{0}(M,Ad\,P)$, we have the relation
$\langle[\eta_{1},\eta],\eta_{2}\rangle=\langle\eta_{1},[\eta,\eta_{2}]\rangle.$ (2.2)
@@
With (2.1) and (2.3), we can also express $\langle-,-\rangle_{\mathcal{C}}$ as an integral by introducing the Hodge star operator for cone forms
@@
Having defined the inner product on $\operatorname{Cone}^{*}(\zeta)(M,Ad\,P)$, we write down the Yang-Mills functional associated with the cone connection form. Given a connection form $A$ of $Ad\,P$ over $M$ and a section $B\in\Omega^{0}(M,Ad\,P)$, we write the cone connection form as $\mathcal{A}=A+\theta B$. Locally, if we write the usual covariant derivative as $d_{A}=d+A$, then the cone covariant derivative is given by $D_{\mathcal{C}}=d+\mathcal{A}=d+(A+\theta B)$ and the cone curvature takes the form
@@ 2.3 Comparison with Yang-Mills solutions
Let $\Sigma$ be a Riemann surface and $A^{\prime}$ a non-flat Yang-Mills connection on $\Sigma$. Let $\zeta$ be the volume form of $\Sigma$ normalized such that the total volume of $\Sigma$ is one. Define $M$ to be the three-dimensional circle bundle $\pi:M\to\Sigma$ with Euler class given by $\zeta$. Since $A^{\prime}$ is Yang-Mills, we can write its curvature as $F_{A^{\prime}}=\zeta\Phi$ such that $d_{A^{\prime}}\Phi=0$. For simplicity, we will also use $\zeta$, $A^{\prime}$ and $\Phi$ to denote their pullbacks on $M$.
@@
With $\zeta$ being a non-degenerate two-form, (2.21) gives the condition $dB=0$.
@@ 2.4 Uniqueness of cone Yang-Mills solutions for a fixed connection
However, it is interesting to ask that given a cone Yang-Mills solution $(A,B)$, how many different $B$’s with $A$ fixed would also be a cone Yang-Mills solution? In the case of $\zeta=0$ and $M$ closed, if $(A,B)$ is a cone Yang-Mills solution, then so are all $(A,B+\Phi)$ with $d_{A}\Phi=0$. For $\zeta\neq 0$, we are able to obtain results in two special cases. First, in two dimensions and suppose $A$ is a Yang-Mills connection, we have the following:
@@
Let $\dim M=2n$. With respect to a compatible metric, $*\,\zeta=*\,\omega=\frac{1}{(n-1)!}\omega^{n-1}$, and $\zeta^{*}\zeta=n$.
@@ 3.2 Three-dimensional solutions from duality relations
conformally scaled by the $e^{2\phi}$ factor in order to match the desired metric form of $ds^{2}_{X}=ds^{2}_{M}+\theta^{2}$ as in (1.5). We note that the four-dimensional Yang-Mills anti-self-dual/self-dual curvature equation is invariant under an overall conformal scaling of the metric. Therefore, for the cone pair $(A,B)$ on $M=\mathbb{R}^{3}-\{0\}$, we can just read off from (3.11)-(3.12) by setting $\mathcal{A}^{r}{}_{s}=(A+\theta B)^{r}{}_{s}=\omega^{r}{}_{s}\,$. We find
@@
We can consider dimensionally reducing the three-dimensional duality condition $*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\mp\,\mathcal{F}_{\mathcal{A}}$ over another circle down to a two-dimensional manifold $N$. For the metric, we will assume $ds_{M}^{2}=ds_{N}^{2}+\chi^{2}$ where $\chi$ is the global connection one-form (i.e. the global angular form) of the circle bundle $M$ over $N$. The dimensionally reduced three-dimensional connection can then be expressed as $A+\chi b$ where $A$ is now a connection form over $N$ and $b\in\Omega^{0}(N,Ad\,P)$. The three-dimensional cone curvature then takes the form
@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
To simplify our consideration, we will additionally assume that $\zeta$ is a non-degenerate two-form. Being $d$-closed and non-degenerate, $\zeta$ is then a symplectic structure and $M$ must be even-dimensional. Additionally, $\zeta$ being non-degenerate leads to two simplifications in considering cone-flat pairs $(A,B)$. First, non-degeneracy ensures that $\zeta$ nowhere vanishes, and hence, the first cone-flat equation $F_{A}=-\zeta B$, implies $B$ is determined by the connection $A$. Second, the second cone-flat equation $d_{A}B=0$ is automatically satisfied and redundant when $\dim M\geq 4$. This is because the first cone-flat equation together with the Bianchi identity imply
@@
Data Availability There is no data associated with this article.
Conflict of Interest The authors have no financial or proprietary interests in any material discussed in this article.